OpenStudy (anonymous):

Solve this equation for x using exponent properties: 3/4x^-2/3=2.4 . Give the answer correct to the nearest thousandth. Please and tanks!

OpenStudy (anonymous):

kinda late but did u get it i was eating lol

OpenStudy (anonymous):

lol no :( not this one or another one

OpenStudy (anonymous):

alryte first move the x to the numerator so it becomes a positive exponent. so you have (3x^2/3) / 4. you can rewrite that as .75 times the cube root of x^2 = 2.4. devide both sides .75, then u have cube root of x^2 equals 3.2. then cube both sides to get rid of the cube root. youll have x^2 equals 32.768. then square root both sides and youll get x

OpenStudy (anonymous):

i hope lol

OpenStudy (anonymous):

loll that actually makes sense!! thank you!! would you mind helping me with the one on chat?

OpenStudy (anonymous):

alryte

OpenStudy (anonymous):

Fossils found in ancient Indian ruins were found to have 34% of carbon-14 remaining in them. How old are the fossils? Show how to determine this using the function in part B. Part B: In(P)/In(1/2)= T/5700

OpenStudy (anonymous):

THANK YOU THANK YOU THANK YOU SO MUCH!

OpenStudy (anonymous):

this precal class?

OpenStudy (anonymous):

nahh algebra 2

OpenStudy (anonymous):

do you have the answer? i got 8871.44

OpenStudy (anonymous):

how??

OpenStudy (anonymous):

just plugged in the .34 for P then multiply both sides by 5700

OpenStudy (anonymous):

but did you use the formula?

OpenStudy (anonymous):

ya ln(.34)/ln(1/2) = T/5700

OpenStudy (anonymous):

check in the back of your book to make sure my answer is right

OpenStudy (anonymous):

it doesnt tell us the answer

OpenStudy (anonymous):

great lol

OpenStudy (anonymous):

lol i know right!? just my luck!

OpenStudy (anonymous):

well thats what i think the answer is good luck!

OpenStudy (anonymous):

lol i tried it again and i got tht! thanks!

OpenStudy (anonymous):

np

OpenStudy (anonymous):

:)